what two whole numbers numbers in a row whose sum is 36

Answer by oberobic(2304)

x+y = 36
y = x+1
.
2x+1 = 36
2x = 35
x = 17.5, which is not a whole number
.
17+18 = 35
18+19 = 37
.
BUT, perhaps there are two consecutive even or odd numbers that total 36?
.
x + y = 36
y = x+2
.
x +x+2 = 36
2x = 34
x = 17
x+2 = 19
.
So 17 and 19 are two consecutive odd numbers that total 36.
.
Done.